Mrs. Hool is planting a garden in her backyard this summer. It must be fenced in to keep out animals. She only has 50 feet of fencing. She would like the garden to be at least 10 feet long and at least 5 feet wide. What is a graph showing the possible dimensions her garden could be?
Name two possible dimensions for her garden.

Answer by ikleyn(52781)

.
The inequalities are

L >= 10 ft    (length)
W >=  5 ft    (width)

2L + 2W = 50 ft   (perimeter)    or, equivalently
 L  + W = 25 ft.


The graph showing possible dimensions of her garden is presented in the Figure below.


The horizontal axis is L. Vertical axis is W.





Plot L + W = 25 (red), L = 10 (green)  and W = 5 (blue)


The solution set is the straight line segment = the hypotenuse of the triangle, shown in red.


The two possible combinations are the highest vertex of the triangle and the rightmost vertex of the triangle.


Try to cipher these points on your own.